OpenStudy (samigupta8):

Let f:R------>[1,infinity) be defined as f(x)=log(√3x^2-4x+k+1 +10).If f(x) is surjective then

OpenStudy (samigupta8):

Root is not over the number 10

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

I made conditions for domain and then I m not getting how shall i proceed

OpenStudy (michele_laino):

I think that we have to rewrite such conditions, in a more useful way

OpenStudy (michele_laino):

for example, we can rewrite those conditions as below: \[\Large \begin{gathered} {x^2} - \frac{4}{3}x + \frac{{k + 1}}{3} \geqslant 0 \hfill \\ \hfill \\ \left( {\sqrt {{x^2} - \frac{4}{3}x + \frac{{k + 1}}{3}} + \frac{{10}}{{\sqrt 3 }}} \right) > 0 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

from such condiions, I get: \[\Large \left\{ \begin{gathered} {\left( {\frac{2}{3}} \right)^2} - \frac{{k + 1}}{3} \leqslant 0 \hfill \\ \hfill \\ {\left( {\frac{2}{3}} \right)^2} - \left( {\frac{{k + 1}}{3} - \frac{{100}}{3}} \right) < 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (samigupta8):

From your conditions i got k>1/3 And k>301/3 which implies that k>301/3... Isn't it?